OpenStudy (anonymous):

The length of an auto racetrack loop is 1.855 miles. On a test run, Steve completes 35 laps in 30 minutes. In miles per hour, what was Steve's average rate of speed during his test run?

OpenStudy (anonymous):

I mean 37.1

OpenStudy (jack1):

hmm, could u show me ur steps plz, i got a different anwser @ctfxcfan1

OpenStudy (anonymous):

I did all the same steps I did in the last one but I may have done something wrong in the mix I'm re doing it right now

OpenStudy (jack1):

ok no worries

OpenStudy (anonymous):

I think I maybe skipping a step or 2 I know it probably has a decimal in the answer A 64.925 B 37.1 C 70 D 129.85

OpenStudy (jack1):

srsly dude, post it up, step by step what ur doin, and we'll go through it together hey

OpenStudy (anonymous):

1.885 times 35=64.925 64.935/30 times 60/1=129.85 I forgot the dividing

OpenStudy (anonymous):

http://www.twiddla.com/1514607

OpenStudy (jack1):

ah ha, that looks better u needed ur time in hours so 30 mins = 1/2 hrs = 0.5 hrs distance u did perfectly so speed = dist / time = 129 ish